Determine the volume, in milliliters, for each of the following:
a.) a 1.9 M NaOH solution prepared from 60.0 mL of a 16 M NaOH solution
Express your answer to two significant figures and include the appropriate units.
b.) a 1.3 % (m/v) CaCl2 solution prepared from 22 mL of a 5.4 % (m/v) CaCl2 solution
Express your answer to two significant figures and include the appropriate units.
c.)a 0.50 M HCl solution prepared from 49 mL of a 8.7 M HCl solution
Express your answer to two significant figures and include the appropriate units.
d.)a 1.15 % (m/v) H2SO4 solution prepared from 10.9 mL of a 19.4 % H2SO4 solution.
Express your answer to three significant figures and include the appropriate units
1
Expert's answer
2018-04-07T08:17:11-0400
a) n(NaOH) = 16 mol/L*0.06L = 0.96 mol
V = 0.96 mol / 1.9 mol/L = 0.505 L = 505 mL
b) m(CaCl2) = 22*0.054 = 1.188 g
V = 1.188 / 0.013 = 91 mL
c) n(HCl) = 8.7 mol/L*0.049 L = 0.43 mol
V = 0.43 mol / 0.5 mol/L = 0.86 L = 860 mL
d) m(H2SO4) = 10.9*0.194 = 2.1146 g
V = 2.1146 / 0.0115 = 184 mL